### v1

# Accelerating Data Chunking in Deduplication Systems using Vector Instructions

Abstract.
Content-defined Chunking (CDC) algorithms dictate the overall space savings that deduplication systems achieve. However, due to their need to scan each file in its entirety, they are slow and often the main performance bottleneck within data deduplication. We present VectorCDC, a method to accelerate hashless CDC algorithms using vector CPU instructions, such as SSE / AVX. Our evaluation shows that VectorCDC is effective on Intel, AMD, ARM, and IBM CPUs, achieving $8.35\times-26.2\times$ higher throughput than existing vector-accelerated techniques without affecting the deduplication space savings.
Content-defined chunking, SSE/AVX instructions, AVX-512, ARM NEON, IBM VSX
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXcopyright: noneccs: Information systems Cloud based storageccs: Networks Cloud computingccs: Information systems Deduplicationccs: Computer systems organization Single instruction, multiple data
1. Introduction
The amount of data generated and stored on the Internet is growing at an exponential rate (statista2024, ), and is expected to exceed 180 zettabytes per year in 2025. Storage capacity alone is not well positioned to handle this data influx, with the total installed storage capacity in 2020 only being 6.7 zettabytes (statista2024, ). Cloud storage providers instead support this data growth using alternatives such as novel storage paradigms (software_defined_storage, ; raid, ), distributed file systems (hadoop, ; ceph, ) and caches (memcached, ; facebook_tao, ), mechanisms such as data deduplication (dedup_intro, ; comprehensive_dedup, ), alongside additional investments in data protection (data_security, ).
Previous studies by Microsoft (primary_dedup, ) and EMC (backup_workload_charz, ) show that a large amount of redundancy exists in the data stored on the cloud, especially in file system backups (primary_dedup, ; dedup_intro, ), virtual machine backups (backup_workload_charz, ; wandeltacomp, ) and shared documents (documentduplication_2011, ). The redundancy ratios in these workloads are significant, ranging from $50\%$ to $75\%$. Mechanisms such as data deduplication (dedup_intro, ) and compression (wandeltacomp, ) are used to conserve storage space by identifying redundant data, eliminating it, and minimizing its storage impact, thereby improving efficiency and reducing costs.
Data deduplication consists of four phases (comprehensive_dedup, ); Data Chunking, Chunk Fingerprinting, Metadata Creation, and Data Storage. Data chunking and chunk fingerprinting are the most computationally intensive (dedup_intro, ; dedupbench, ) of these. While chunk fingerprinting has received significant optimization attention, with faster hashing algorithms (tunable_encrypted_dedup, ; scalable_incremental_checkpt, ) and GPUs (gpu_dedup, ; storegpu, ), data chunking optimizations have not kept up (§3.1).
In the data chunking phase, the incoming data is divided into small chunks, typically of size 1641-64 KB. Numerous data chunking algorithms exist in current literature (ae, ; fastcdc, ; gear_hash, ; lbfs, ; ram, ; tttd, ) and can be broadly classified into hash-based and hashless algorithms (dedupbench, ). As chunking occurs whenever new data is uploaded, this phase is on the critical path, and directly impacts system performance.
Previous efforts explored accelerating chunking through using vector instructions. SS-CDC (sscdc, ) uses vector instructions to accelerate hash-based data chunking algorithms, such as Rabin-Karp chunking (lbfs, ) and Gear-based chunking (gear_hash, ). Unfortunately, this approach only leads to modest improvements in chunking throughput, up to $3.13\times$, as shown in §3.2. Parallelizing hash-based chunking using vector instructions is complicated because these algorithms use the rolling hash of a sliding window of bytes to detect boundaries, inherently creating a computational dependency between adjacent bytes. Consequently, SS-CDC processes different regions of the data in parallel. This approach relies on slow scatter / gather vector instructions, limiting its performance (§3.2).
We posit that hashless chunking algorithms are better candidates for vector acceleration. Although they achieve slightly lower space savings compared to their hash-based counterparts (§6), they are up $2\times$ faster and use simple mathematical operations (e.g., finding a maximum value) that can be accelerated more efficiently using vector instructions.
We analyzed state-of-the-art hashless algorithms to understand their design and identify opportunities to leverage vector instructions. We identified that all state-of-the-art hashless algorithms consist of two processing patterns. The first pattern involves finding local minima or maxima in a data region, which we call Extreme Byte Search, and the second pattern involves scanning a range of bytes to find values that are greater or lesser than a target value, which we call Range Scan. We found that, unlike rolling hash functions, these patterns can be efficiently accelerated using vector instructions.
Using these insights, we present VectorCDC, a technique for accelerating hashless chunking algorithms using vector instructions. VectorCDC uses a novel design to accelerate the two aforementioned patterns. We accelerate the extreme byte searches with a novel tree-based search that divides a region of bytes into multiple sub-regions, processes each region using vector instructions, and uses a tree-based approach to combine their results. We accelerate range scans with packed scanning, which packs multiple adjacent bytes into vector registers and compares them using a single vector operation.
We implemented VectorCDC using five different vector instruction sets: SSE-128, AVX-256, and AVX-512 on Intel / AMD CPUs; NEON-128 on ARM CPUs; and VSX-128 on IBM Power CPUs. We used VectorCDC to accelerate three state-of-the-art hashless chunking algorithms. We implemented VRAM, VAE, and VMAXP, which are accelerated versions of the RAM (ram, ), AE (ae, ), and MAXP (maxp, ) algorithms, respectively. We compared the performance of our accelerated algorithms against state-of-the-art hash-based algorithms, hashless algorithms, and SS-CDC (sscdc, ) accelerated algorithms using 10 diverse datasets.
Our evaluation (§6) shows that VectorCDC achieves $8.35\times$$26.2\times$ higher chunking throughput than SS-CDC’s approach, without affecting the space savings achieved by hashless algorithms. VectorCDC-based algorithms achieve up to $5.51\times$$17.6\times$ higher throughput compared to their unaccelerated counterparts. Furthermore, algorithms accelerated using VectorCDC retain their performance advantage with all five vector instruction sets. We have made our code publicly available by integrating it with DedupBench [^1] https://github.com/UWASL/dedup-bench (dedupbench, ). Due to the large sizes of our datasets (§6), we were unable to release them all. We have instead provided detailed descriptions to facilitate their recreation for future experiments, similar to previous literature (fastcdc, ; ae, ; ram, ). In addition, to facilitate reproducing our results, we have publicly released one of our datasets (DEB) on Kaggle [^2] https://www.kaggle.com/datasets/sreeharshau/vm-deb-fast25 (deb_dataset, ).
The rest of this paper is organized as follows: we discuss relevant background about deduplication and vector instructions in Section 2. Section 3 motivates our work by discussing deduplication performance bottlenecks and the inefficiencies encountered by previous work when accelerating hash-based CDC algorithms. Section 4 outlines VectorCDC’s design while Section 6 details our evaluation efforts on diverse datasets. We discuss related work in Section 7 and conclude our paper in Section 8.
2. Background
Data deduplication consists of four phases (comprehensive_dedup, ):
  • Data Chunking: Data is divided into small chunks typically of size 1641-64KB using a chunking algorithm. All chunking algorithms have configurable parameters that control the average size of generated chunks.
  • Fingerprinting and Comparison: Chunks are hashed using a collision-resistant hashing algorithm such as MurmurHash3 (tunable_encrypted_dedup, ) or SHA-256 (sha256, ) to generate fingerprints. Fingerprints are compared against those previously seen to identify duplicate chunks.
  • Metadata Creation: Metadata, i.e., file recipes required to reconstruct the original data from stored chunks, are created.
  • Metadata and Chunk Storage: Non-duplicate chunks and recipes are saved on the storage medium. Fingerprints are stored on the fingerprint database and cached in an in-memory index.
Data chunking and fingerprinting are typically the most computationally intensive phases in deduplication (dedup_intro, ). While fingerprinting has been accelerated up to $53\times$ using GPUs (gpu_dedup, ; storegpu, ) and faster hashing algorithms (tunable_encrypted_dedup, ; persistent_memory_dedup, ), data chunking acceleration has only received limited attention and is currently the performance bottleneck (§3.1).
2.1. Data Chunking
Data chunking algorithms can generate fixed-size or variable-sized chunks. Dividing the data into fixed-size chunks is fast, but results in poor space savings on most datasets. This is due to the byte-shifting problem (lbfs, ), where adding a single byte causes all subsequent chunks to appear different, despite the data stream largely being unchanged. Thus, while traditional backup systems such as Venti (venti, ) and OceanStore (oceanstore, ) use fixed-size chunks, modern deduplication systems employ content-defined chunking (CDC) algorithms (lbfs, ) to generate variable-sized chunks.
Chunk boundaries in CDC algorithms are derived from the data itself, i.e., they are content-defined. These boundaries are chosen such that most byte shifts cause them to shift by the corresponding amount, leaving subsequent chunks unaffected and preserving the ability to detect duplication. Numerous CDC algorithms have been proposed in previous literature (ae, ; fastcdc, ; gear_hash, ; lbfs, ; maxp, ; ram, ; tttd, ) and can be broadly classified into hash-based and hashless algorithms (dedupbench, ).
(a) Sliding window and rolling hash
(b) Minimum chunk sizes and sub-minimum skipping
Figure 1. Hash-based chunking algorithms
2.1.1. Hash-based algorithms.
These algorithms (fastcdc, ; gear_hash, ; lbfs, ; tttd, ) slide a fixed-size window over the data. When the hash value of the window’s contents matches a target mask, they insert a chunk boundary, creating a new data chunk lying between the current and previous chunk boundaries. Note that these hash-based CDC algorithms are only used during the Data Chunking phase and do not affect the Fingerprinting and Comparison phase.
Algorithms such as Rabin-Karp chunking (lbfs, ) and CRC (sscdc, ) slide a window over the source data and compute the hash of the window’s contents using rolling hash algorithms. Figure 1(a) shows an example of data chunking with such algorithms. In the Rabin-Karp chunking algorithm (lbfs, ), a chunk boundary is declared when the lower k bits of the sliding window’s hash value equals zero. If the current window’s hash value does not meet this condition, the window is slid by a byte. To minimize the overhead of recomputing the hash value, the new value is calculated as a function of the old hash value, the incoming byte, and the outgoing byte (Figure 1(a)), i.e., a rolling hash. This creates a dependency between adjacent bytes, complicating acceleration efforts with SIMD instructions (§3.2). Additionally, despite the development of more lightweight rolling hash algorithms such as CRC (sscdc, ) and Gear-Hash (gear_hash, ), hash-based chunking remains computationally expensive (§6.2).
Some hash-based algorithms like TTTD (tttd, ) and FastCDC (fastcdc, ) use minimum and maximum values to limit the chunk sizes. To improve chunking throughput, these algorithms skip scanning data lying before the minimum chunk size at the beginning of each chunk. Figure 1(b) shows an example of such algorithms with the sub-minimum regions highlighted using a dashed pattern. To offset the impact of skipping the sub-minimum regions and tighten chunk size distributions around the average, FastCDC (fastcdc, ) uses dynamically changing masks, i.e., relaxes the boundary detection condition by reducing k when required. TTTD (tttd, ) uses two different boundary masks to do the same.
2.1.2. Hashless algorithms.
[…]
AE. Figure 2(a) shows an example chunk generated by the Asymmetric Extremum (AE) (ae, ) algorithm. AE has two modes of operation: AE-Min and AE-Max, depending on whether it uses local minima or maxima; Figure 2(a) shows AE-Max. In each chunk, AE-Max tries to find a target byte that is greater than all the bytes before it. Once the target byte is identified, AE-Max scans a fixed-size window of bytes after the target to identify the maximum-valued byte among them. If the target byte is greater than this maximum-valued byte, it inserts a chunk boundary after the fixed-size window, as shown in Figure 2(a).
Similarly, AE-Min tries to find a target byte that is less than all the bytes before it. When such a byte is identified, AE-Min scans a fixed-size window of bytes after the target to identify the minimum-valued byte within. If the target byte is lesser than this minimum-valued byte, it inserts a chunk boundary after the fixed-size window.
MAXP. Figure 2(b) shows an example chunk generated by MAXP (maxp, ). MAXP identifies target bytes in the data stream that are local maxima, i.e., they are greater than a fixed number of bytes before and after them. When such target bytes are found, chunk boundaries are inserted at their locations, as shown in the figure. Note that MAXP’s window sizes are typically 70-80% smaller than AE (ae, ) and RAM (ram, ) to generate the same target average chunk size.
(a) AE-Max
(b) MAXP
(c) RAM
Figure 2. Hashless chunking algorithms
MAXP works by sliding two fixed-size windows over the data, tracking the maximum values from both windows. These windows are located one byte apart, as shown in Figure 2(b), and the byte between them is the target byte. When the target byte’s value is greater than the maximum value from both windows, a chunk boundary is inserted as the target byte is a local maximum.
RAM. Figure 2(c) shows an example chunk generated by the Rapid Asymmetric Maximum (RAM) (ram, ) algorithm. RAM begins by scanning a fixed-size window at the beginning of each chunk to find the maximum valued byte (F1 in the figure). It then begins scanning at the first byte outside the window, serially comparing these bytes against this maximum value. A chunk boundary is inserted when the first byte that exceeds or equals the maximum is found, e.g., F3 in Figure 2(c).
As they do not possess explicit dependencies between adjacent bytes, we argue that hashless algorithms are better candidates for SIMD acceleration efforts.
2.2. Deduplication Metrics
Previous literature (understanding_dedup_ratios, ; dedupbench, ; lbfs, ; fastcdc, ) outlines three important metrics for deduplication systems: Space Savings, Chunk Size Distribution, and Chunking Throughput. We describe these in detail in this section.
2.2.1. Space savings.
Space savings (understanding_dedup_ratios, ; dedupbench, ) is one of the primary metrics used to evaluate deduplication systems in production. It represents the overall disk space conserved by using the deduplication system, i.e. size of data stored on disk after deduplication. The space savings achieved are largely dictated by the choice of the data chunking algorithm and its associated parameters. It is defined as:
(1) $Space~Savings~(\%)=\frac{Original~Data~Size-Deduplicated~Data~Size}{Original~Data~Size}\times 100$
2.2.2. Chunking throughput.
Chunking throughput is defined as the speed at which the deduplication system divides incoming data into chunks. As CDC algorithms are content-dependent, they need to scan every byte of an incoming data stream before making content-defined boundary decisions. Their speed depends on their computational complexity. Hash-based algorithms involve the use of expensive rolling-hash algorithms to determine chunk boundaries (§2.1), typically resulting in lower throughputs than their hashless counterparts. Chunking throughput is defined as:
(2) $Chunking~Throughput=\frac{Original~Data~Size}{Time~taken~to~generate~all~chunk~boundaries}$
2.2.3. Chunk size distribution.
Content-defined chunking (CDC) algorithms generate variable-sized chunks of a target average chunk size. They try to ensure that the sizes of generated chunks are as close to the target average as possible. However, due to underlying algorithmic characteristics, each algorithm has a unique chunk size distribution pattern. For instance, FastCDC (fastcdc, ) exhibits two distinct smooth distributions, changing its pattern at the average chunk size. This is because it switches masks past the target average size and relaxes boundary conditions. On the other hand, algorithms such as TTTD (tttd, ) exhibit a smooth distribution between their minimum and maximum specified chunk sizes. Chunk size distributions are typically represented using cumulative distribution function (CDF) plots.
Space savings are inversely proportional to the target average chunk size, i.e., the greater the average chunk size, the lower the space savings achieved in general (primary_dedup, ). This is because the probability of finding duplicate chunks is higher at smaller chunk sizes. The degree of space savings degradation with increasing chunk size depends on algorithmic and dataset characteristics.
All chunks generated by CDC algorithms are subsequently hashed using a collision-resistant algorithm (tunable_encrypted_dedup, ) to generate fingerprints, as described above. The set of unique fingerprints observed thus far is stored in a fingerprint database. New incoming chunks are hashed, and their fingerprints are compared against this database to detect duplicates. Thus, smaller and more numerous chunks result in a larger database; specifically, the fingerprint database size is inversely proportional to the chosen average chunk size. A large number of small chunks can negatively impact system throughput, both due to the increased fingerprint database size and the random data accesses caused by these chunks. Thus, CDC algorithms in production typically target average chunk sizes between 22KB–6464KB.
2.3. Vector Instructions
Vector instructions (vector_instruction_sets, ) are special Single-Instruction Multiple-Data (SIMD) instructions supported by most modern processors. They rely on special vector registers for their operations. These registers come in multiple sizes; depending on the width of the vector register they use, vector instructions can be classified into different families (vector_instruction_sets, ). The most common vector register sizes are 128 bits, 256 bits, and 512 bits, i.e., 16, 32, and 64 bytes wide.
Vector instructions support the execution of an operation on multiple pieces of data by packing them into vector registers; for instance, eight 16-bit values $a$-$h$ can be densely packed into a 128-bit vector register $V_{1}$. To add $a$-$h$ to eight other values $i$-$p$, we can pack $i$-$p$ into another register $V_{2}$. We can now add them pairwise with a single vector addition operation VADD ($V_{1},V_{2}$) using $V_{1}$ and $V_{2}$ as operands, instead of eight separate integer arithmetic operations.
Vector instructions support various arithmetic operations (intelIntrinsicsGuide, ), including pairwise addition, subtraction, multiplication, and maximum/minimum on packed values. Additionally, they support logical operations such as bitwise AND (&) and bitwise OR (|). They have been previously used to accelerate matrix multiplication (avx_matrixmul, ), sorting (avx_quicksort, ), multimedia applications (avx_multimedia, ), fluid simulations (avx_fluidmech, ), hash tables (hashtables_vector, ), and relational databases (fused_table_scan, ). The supported vector instruction types and their relative performance vary across CPU manufacturers.
2.3.1. Intel and AMD
Vector instructions on x86 platforms can be classified into three families: SSE-128, AVX-256, and AVX-512 (vector_instruction_sets, ). SSE-128 instructions use 128-bit registers and have been supported by Intel and AMD processors since 1999 (intelSIMDInstructions, ) and 2003 (AMD_SSE128, ), respectively. AVX-256 instructions were introduced by Intel and AMD in 2011 (intelSIMDInstructions, ), and use 256-bit registers. Finally, only a handful of the newest Intel (intelXeonPhi, ) and AMD (amdzen4_wikichip, ) processors, which have 512-bit wide vector registers, support AVX-512 instructions.
2.3.2. ARM
ARM processors have supported NEON-128 instructions, an equivalent to SSE-128, since 2011 (armNeonOverview, ). Modern ARM processors also support 256-bit and higher vector widths with the SVE / SVE2 instruction sets, available since 2021 (arm_sve, ). These two instruction sets differ in the kinds of instructions supported. For instance, NEON-128 does not support native VMASK operations, which are used to create integer masks by extracting one out of every k bits in a vector register, while SVE / SVE2 does. This can lead to performance differences in applications that need the VMASK operation (hashtables_vector, ).
2.3.3. IBM Power
IBM’s Power (ibm_power8, ) architecture supports AltiVec / VSX-128 vector instructions (altivec_instructions, ), an equivalent to SSE-128, since the 1990s. This instruction set supports equivalents for most SSE instructions but lacks VMASK support.
3. Motivation
3.1. Performance bottlenecks in data deduplication
Figure 3. Time taken for Data Chunking vs Fingerprinting while deduplicating randomized data with FastCDC (fastcdc, ) and an 8 KB average chunk size.
Although both data chunking and fingerprinting have traditionally been considered the main bottlenecks in deduplication (dedup_intro, ), this has changed with the advent of new hashing algorithms/acceleration methods for fingerprinting. Figure 3 is a stacked bar plot showing the time taken by the data chunking and fingerprinting phases during deduplication. For fingerprinting, we use five different collision-resistant hashing algorithms (md5, ; sha1, ; sha256_512, ; appleby2008murmurhash, ; hashing_survey_xxhash, ). For data chunking, we use FastCDC (fastcdc, ), the fastest unaccelerated chunking algorithm (§6.2), with an 8 KB average chunk size. We use 30 GB of randomized data and an Intel Emerald Rapids machine described in §6 for this experiment.
Figure 3 shows that fingerprinting and data chunking take nearly equal time with traditional collision-resistant hashing algorithms, such as SHA1 (sha1, ) and SHA-256 (sha256_512, ). Fingerprinting takes longer than data chunking with other algorithms, such as MD5 (md5, ) and SHA-512 (sha256_512, ). This indicates that both phases used to be the performance bottlenecks in deduplication, aligning with previous literature (dedup_intro, ).
However, recent research (tunable_encrypted_dedup, ; scalable_incremental_checkpt, ) has introduced faster hashing algorithms such as MurmurHash3 (appleby2008murmurhash, ) and xxHash3 (xxhashWebsite, ). Figure 3 shows that fingerprinting takes significantly lower time than chunking with these new hashing algorithms, as they are $10\times-15\times$ faster than their counterparts on CPUs. Using GPUs can further accelerate fingerprinting speeds by up to $53\times$ (gpu_dedup, ; storegpu, ). Note that the choice of fingerprinting algorithm does not affect deduplication space savings, as MurmurHash3 and xxHash3 have similar collision resistance properties to other algorithms (appleby2016smhasher, ).
Thus, as a result of its computationally intensive nature and position on the critical path, data chunking is the main performance bottleneck in the deduplication pipeline and needs further optimization.
3.2. Accelerating hash-based algorithms with vector instructions
Figure 4. SS-CDC (sscdc, ): Accelerating the rolling hash phase
To address the data chunking bottleneck, SS-CDC (sscdc, ) proposed using AVX-512 instructions to accelerate hash-based CDC algorithms. They decouple the rolling hash and boundary detection phases, running the rolling hash on the entire source data to identify boundary candidates in the first phase, and determining boundaries sequentially in the second. This allows both stages to be independently accelerated with AVX instructions.
Figure 4 shows how SS-CDC (sscdc, ) accelerates the first rolling hash phase of hash-based CDC algorithms. SS-CDC uses AVX-512 registers to create multiple rolling-heads, i.e., calculating the rolling hash on bytes from multiple regions of the file independently and in parallel. Each rolling head maintains its own hash value and independently calculates the contributions of incoming and outgoing bytes.
To use vector instructions, they first collect the outgoing bytes for each head into a vector register $V_{1}$. Similarly, they collect all the incoming bytes into another vector register $V_{2}$. The hash values for each head are stored in a separate register $V_{3}$. SS-CDC removes the contributions of all outgoing bytes from the hash values with a single vector operation and adds the contributions of all incoming bytes with another. Whenever any of the hash values match the boundary condition (such as the lower k bits being equal to zero in Rabin-Karp chunking (lbfs, )), they mark the current position as a boundary candidate in a separate bitmap. In the second phase, they scan the bitmap using vector instructions to determine the actual boundaries among all candidates, taking into account the minimum and maximum chunk sizes.
This approach introduces two problems. First, many hash-based algorithms, such as TTTD and FastCDC, skip scanning data up to the minimum chunk size to improve throughput (§2.1). Decoupling the rolling hash and boundary detection phases causes the rolling hash to be run on the entire incoming data stream, nullifying these optimizations.
Second, to load incoming and outgoing bytes from different regions in the file, SS-CDC (sscdc, ) uses AVX gather instructions. To populate the candidate bitmap when boundary candidates are discovered, they use scatter instructions. These scatter and gather instructions are slow (scatter_gather_perf, ), limiting performance gains. Finally, scatter instructions are only available on processors supporting certain instruction sets (intelIntrinsicsGuide, ), limiting SS-CDC’s usage to a handful of the newest Intel and AMD processors (§2.3).
Figure 5. SS-CDC (sscdc, ) throughputs on randomized data with AVX-512 instructions
Figure 5 shows the chunking throughput obtained by running SS-CDC accelerated versions of CRC (SS-CRC) and Gear-based chunking (SS-Gear) (sscdc, ) against their native unaccelerated counterparts. This experiment used randomized data, an Intel Emerald Rapids machine described in §6 and AVX-512 instructions. We ran each algorithm with chunk sizes of 4164-16 KB. SS-CRC achieves 1.2 GB/s, a speedup of $2.58\times$ over CRC. Similarly, SS-Gear achieves 3.3 GB/s, a speedup of $3.13\times$ over Gear. These small speedups result from the challenges associated with hash-based algorithms that are described above.
Hashless algorithms do not possess explicit dependencies between adjacent bytes. They treat each byte as an independent value and use maximum / minimum values from data regions to determine chunk boundaries. VectorCDC chooses hashless algorithms over their hash-based counterparts as they are better candidates for SIMD acceleration.
4. VectorCDC Design
Figure 6. Accelerating Extreme Byte Search
Hashless CDC algorithms such as AE (ae, ), RAM (ram, ), and MAXP (maxp, ) slide windows over the source data to determine chunk boundaries. We identify two common processing patterns across all hashless CDC algorithms: the Extreme Byte Search and Range Scan. We accelerate each of these patterns using different vector-based techniques, which are discussed in detail below.
While we use the AVX-512 instruction set as an example to describe our acceleration techniques in this section, they can be implemented on any CPU with a vector instruction set supporting VMAX, VCMP and VMASK operations. SSE-128 and AVX-256 instruction sets (intelIntrinsicsGuide, ) fall under this umbrella, as do ARM and IBM processors with NEON-128 (armNeonOverview, ) and AltiVec / VSX-128 (ibm_power8, ) instructions, respectively. Thus, VectorCDC is compatible with a wide range of processors, unlike SS-CDC (sscdc, ), which relies on scatter instructions only available in AVX-512 instruction sets. Finally, while other minima/maxima-based hashless algorithms can also be accelerated using VectorCDC, their native versions are slower (ae, ; ram, ; maxp, ) than AE, RAM, and MAXP and have been omitted from the rest of our paper.
4.1. Tree-based Extreme Byte Search
Hashless CDC algorithms such as AE (ae, ), RAM (ram, ), and MAXP (maxp, ) all consist of a subsequence that identifies the extreme byte (maximum/minimum) in a fixed-size window. The size of this window depends upon the expected average chunk size and can be as large as 484-8KB. As this subsequence may need to be performed more than once per chunk, we propose accelerating it using a novel tree-based search approach. Let us consider the search for a maximum value using AVX-512 instructions (Figure 6). Note that the same method can be used with other vector instruction sets as well as to find minimum values.
We first divide the fixed-size window into smaller sub-regions, loading all the bytes into AVX-compatible m512i variables in Step 1. We load these bytes in a packed fashion i.e. each m512i variable contains 64 adjacent bytes. We then use vector mm512_max instructions to find the pairwise maximum among packed byte pairs (Step 2). For instance, among bytes ”E1” and ”21”, byte value ”E1” is the maximum. The resulting pairwise maximums are packed into a destination variable ($V_{5}$ in the figure).
Step 3 compares these resulting variables $V_{5}$ and $V_{6}$ from Step 2 using mm512_max instructions to find the pairwise maximums. We repeat this process, building a tree of m512i variables until we are left with a single variable $V_{7}$ containing the maximum-valued 64 bytes from across the entire region. We scan these bytes sequentially in Step K to determine the maximum valued byte.
4.2. Packed Scanning for Range Scans
Figure 7. Accelerating Range Scan
Hashless CDC algorithms also consist of a range scan subsequence, where bytes are serially compared against a target value. We propose to accelerate this scanning process using vector instructions. Let us consider a case where we compare bytes sequentially to see if they are greater than or equal to a target value (such as in RAM (ram, )). Figure 7 shows our proposal to accelerate this using packed scanning with AVX-512 instructions. Note that the same methods are applicable for other vector instruction sets as well.
We first load the target value (”F4” in Figure 7) into an AVX-compatible m512i variable $V_{1}$. We then pack 64 adjacent bytes from the scan region into another m512i variable $V_{2}$. We compare these 2 registers using mm512_cmpge vector compare instructions, which generate a 64-bit integer mask containing the comparison results. If this mask has a value greater than 0, a chunk boundary exists within the scanned 64 bytes. Its exact position is determined using the mask value. If the mask equals 0, no boundary exists within the scanned region and we proceed with loading the next 64 bytes into $V_{2}$ to repeat the process.
Range Scans can be run with one of five comparators: Greater-Than (GT), Lesser-Than (LT), Greater-Than or Equals (GEQ), Less-Than or Equals (LEQ), and exactly Equals (EQ). Each of these comparators uses a different vector compare instruction; for instance, the GEQ comparator uses mm512_cmpge instructions while the LEQ comparator uses mm512_cmple instructions. The same comparators also use different comparison instructions in different instruction sets; for instance, the GEQ comparator uses mm512_cmpge instructions with the AVX-512 instruction set while it uses mm256_cmpge with AVX-256.
It is worth noting that our packed scanning approach is compatible with sub-minimum skipping. Unlike SS-CDC’s approach, chunk boundary detection and insertion can both occur in Range Scans, i.e., whenever a chunk boundary is detected, the next minimum_chunk_size bytes can be skipped.
4.3. Putting it together: AE-Max, AE-Min, MAXP, and RAM
RAM (ram, ) first scans a fixed-size window at the beginning of the chunk to find a maximum value (Figure 2(c)). After this, it inserts a chunk boundary at the first byte outside the window, which is at least as large as the maximum valued byte (§2.1). With VectorCDC, we accelerate RAM as a combination of an Extreme Byte Search to find a maximum value, followed by a Range Scan with the GEQ comparator that compares this maximum value against bytes until a chunk boundary is found.
AE-Max (ae, ) scans for a byte larger than all the bytes before it i.e., a target byte (Figure 2(a)). Once found, a fixed-size window after this byte is scanned to determine the maximum valued byte within. If the target byte is larger than the maximum valued byte, a chunk boundary is inserted; otherwise, scanning continues for a new target byte (§2.1). With VectorCDC, we accelerate AE-Max as a combination of multiple Range Scans with the GT comparator to find target bytes, each followed by a single Extreme Byte Search for a maximum value.
AE-Min (ae, ) scans for a byte with lesser value than all those before it (§2.1). Once found, a fixed-size window after this byte is scanned to determine the minimum value within. If the target byte has a lesser value than the minimum valued byte, a chunk boundary is inserted; otherwise, scanning continues for a new target byte. Similar to AE-Max, we accelerate AE-Min as a combination of multiple Range Scans with the LT comparator to find target bytes, each followed by a single Extreme Byte Search for a minimum value.
Finally, MAXP (maxp, ) scans for a target local maxima that is exactly centered between two fixed-size windows (Figure 2(b)). A chunk boundary is inserted right after such a byte is found (§2.1). Thus, each chunk in MAXP can be represented as a combination of multiple Range Scans with the GT comparator, each followed by two Extreme Byte Searches for maximum values.
5. Implementation
We accelerate AE (ae, ), MAXP (maxp, ), and RAM (ram, ) using VectorCDC with 3000 lines of C++ code. We implemented SSE-128, AVX-256, AVX-512, NEON-128, and VSX-128 versions of all algorithms. We also implemented Extreme Byte Searches for minima and maxima, as well as Range Scan functionalities with the GT, GEQ, LT, LEQ, and EQ comparators on all five vector instruction sets. We have made our code publicly available with DedupBench [^3] https://github.com/UWASL/dedup-bench (dedupbench, ).
Note that while ARM processors support VCMP and VMAX operations, they lack native support for VMASK instructions, which are used during range scans to generate a single mask containing the comparison results. This is a common issue encountered by ARM developers trying to port x86 code (armPortingVector, ). We chose an efficient alternative implementation (armPortingVector, ) to work around the lack of native VMASK support. However, this alternative implementation uses multiple slow NEON-128 instructions, such as vshrn and vreinterpretq, as opposed to a single x86 mm_movemask instruction. As shown in §6.4, this causes accelerated algorithms to achieve lower speedups on ARM CPUs compared to Intel and AMD.
IBM processors also support VCMP and VMAX operations, but lack native VMASK support. However, the same functionality can be achieved using one vec_bperm and two vec_extract instructions. As these instructions are relatively inexpensive, they are an efficient alternative to VMASK. As shown in §6.4, this allows IBM processors to achieve speedups equivalent to or greater than Intel and AMD processors when using VectorCDC.
6. Evaluation
In this section, we evaluate VectorCDC against the state-of-the-art CDC algorithms.
CPU / CPU Family SSE-128 AVX-256 AVX-512 NEON-128 VSX-128
Intel Emerald Rapids - -
Intel Skylake - -
AMD EPYC Rome - - -
ARM v8 Atlas - - - -
IBM Power 8 - - - -
Table 1. Vector instruction sets supported by the different machines in our testbed.
Testbed. We run all our experiments using machines from the Cloudlab (cloudlab_paper, ) platform. We pick five machines with diverse vector instruction set support; Table 1 shows the vector instruction sets supported by each machine. The details of each machine are as follows:
  • Intel Emerald Rapids: We use a c6620 machine from CloudLab Utah, which has a 28-core Intel Xeon Gold 5512U with hyperthreading at 2.1 GHz, 128 GB of RAM, and one Intel NIC each of 25 GBps and 100 GBps. It supports the SSE-128, AVX-256, and AVX-512 vector instruction sets.
  • Intel Skylake: We use a c240g5 machine from CloudLab Wisconsin, which has two 10-core Intel Xeon Silver 4114 CPUs with hyperthreading at 2.2 GHz, 192 GB of RAM, one Mellanox 25 GBps NIC, and one onboard Intel 1 GBps NIC. It supports the SSE-128, AVX-256, and AVX-512 vector instruction sets.
  • AMD EPYC Rome: We use a c6525-25g machine from CloudLab Utah, which has a 16-core AMD 7302P CPU with hyperthreading at 3.0 GHz, 128 GB of RAM, and two Mellanox 25 GBps NICs. It supports the SSE-128 and AVX-256 vector instruction sets.
  • ARM v8 Atlas: We use a m400 machine from CloudLab Utah, which has an 8-core ARM Cortex A-57 CPU at 2.4 GHz, 64 GB of RAM, and a 10 GBps Mellanox NIC. It supports the NEON-128 vector instruction set.
  • IBM Power 8: We use an ibm8335 machine from CloudLab Clemson, which has dual 10-core IBM Power8NVL CPUs at 2.86 GHz with 8 hardware threads per core, 256 GB of RAM, and a 10 GBps Broadcom Xtreme II NIC. It supports the VSX-128 vector instruction set.
While some ARM CPUs released after 2022 support higher vector widths with SVE instructions (§2.3), we could not obtain such a machine for our experiments. Note that all our runs are on the Intel Emerald Rapids machine unless otherwise specified. Our throughput results are the averages of 5 runs, and the standard deviation was less than $5\%$.
Alternatives. We evaluate the following hash-based CDC algorithms:
  • CRC: Native (unaccelerated) version of the CRC-32 chunking algorithm from SS-CDC (sscdc, ).
  • FCDC: Native version of FastCDC (fastcdc, ).
  • Gear: Native version of the Gear-hash based chunking algorithm (gear_hash, ).
  • RC: Rabin’s chunking algorithm from LBFS (lbfs, ).
  • SS-CRC: AVX-512 version of CRC accelerated using SS-CDC (sscdc, ).
  • SS-Gear: AVX-512 version of Gear accelerated using SS-CDC (sscdc, ).
  • TTTD: Two-Threshold Two-Divisor algorithm (tttd, ).
We also evaluate the following hashless CDC algorithms:
  • AE: Native version of the Asymmetric Extremum algorithm (ae, ). We evaluate both AE-Max and AE-Min.
  • MAXP: Native version of the MAXP algorithm (maxp, ).
  • RAM: The native Rapid Asymmetric Maximum (ram, ) algorithm.
  • VAE: Accelerated versions of AE-Max and AE-Min with VectorCDC.
  • VMAXP: Accelerated versions of MAXP with VectorCDC.
  • VRAM: Accelerated versions of RAM with VectorCDC.
Note that for each hashless algorithm accelerated with VectorCDC, we evaluate their SSE-128, AVX-256, AVX-512, NEON-128, and VSX-128 versions on supporting CPU platforms from our testbed (Table 1).
Dataset Size Files Dataset Information XC Median CDC
DEB 40 GB 65
Debian (debian_org, ) VM Images obtained from
the VMware Marketplace (vmware, )
18.98% 34.64%
DEV 230 GB 100 Nightly backups of a Rust (github_rust, ) build server 83.17% 98.05%
FLOW 8 GB 630341 Source files for 25 versions of TensorFlow (tensorflow, ) 90.69% 91.98%
KUBE 1.5 GB 117344 Source files for 5 versions of Kubernetes (kubernetes, ) 64.52% 69.42%
LNX 65 GB 160 Linux kernel distributions (kernel_linux, ) in TAR format (gnu_tar, ) 19.87% 45.62%
MAPS 981 GB 15
OpenStreetMap (openstreetmap, ) backups of Canada
extracted using GeoFabrik (geofabrik, )
0.10% 68.57%
NEWS 478 GB 47
Complete snapshots of a news website across
47 consecutive days in TAR (gnu_tar, ) format
38.95% 73.80%
RDS 122 GB 100 Redis (redis, ) snapshots between redis-benchmark runs 33.54% 92.94%
TPCC 106 GB 25 25 snapshots of a MySQL (mysql, ) VM running TPC-C (TPCCOverview, ) 37.39% 86.64%
WIKI 1 GB 3134
Snapshots of the largest Wikipedia article (wikipediaListFilms, ) across
multiple days, chosen for extreme versioning.
1.31% 72.37%
Table 2. Dataset Information. Note that XC represents the space savings achieved by fixed-size chunking with 8KB chunks while Median CDC is the median space savings achieved by CDC algorithms with an 8KB average chunk size.
Datasets. We use 10 diverse datasets to evaluate VectorCDC; Table 2 shows their details. The datasets represent diverse workloads such as VM backups, database and map backups, web snapshots, and source code. Some datasets, such as FLOW and WIKI, are similar to those used by previous studies (predict_dedup_alberta_2022, ). We have publicly released the DEB dataset  [^4] https://www.kaggle.com/datasets/sreeharshau/vm-deb-fast25 (deb_dataset, ).
We note that the selected datasets have diverse characteristics. They have varying sizes, ranging from $1$ GB for WIKI to $981$ GB for MAPS. They have different file counts; datasets such as MAPS and NEWS consist of a few large files, while others, such as FLOW and KUBE, consist of a large number of small files. We include files with varying formats, such as TAR (gnu_tar, ), VMDK / OVA (ovf_format, ), OSM (openstreetmapFileFormats, ), text files, and binary files across these datasets for comprehensive coverage.
Finally, Table 2 shows the space savings achieved by using fixed-size chunking (XC) and the median of those achieved by CDC algorithms (Median CDC) on these datasets with 8KB chunks. By comparing XC against Median CDC, we note that the datasets possess varying degrees of byte-shifting. For instance, XC achieves a space savings of 37.39% on the TPCC dataset while CDC algorithms achieve a median of 86.68%. On the other hand, XC achieves 90.69% on FLOW while CDC algorithms achieve a median of 91.98%. Two of our datasets, MAPS and WIKI, possess a large degree of byte-shifting, causing an extreme difference in space savings between XC and the median achieved by CDC algorithms.
(a) DEB
(b) DEV
(c) FLOW
(d) KUBE
(e) LNX
(f) NEWS
(g) RDS
(h) TPCC
(i) WIKI
(j) MAPS
Figure 8. Space Savings with 8KB chunks. Note that the legend entries are in the same order as the plot bars.
Metrics. We evaluate the space savings, chunk size distribution, and chunking throughput achieved by each alternative on all the described datasets. In addition, we examine VectorCDC’s performance across different vector instruction sets.
6.1. Space Savings and Chunk Size Distributions
Figures 8(a) - 8(j) show the space savings achieved by all alternatives with 8KB chunks across datasets. We omit the results for other chunk sizes as the trends were similar.
6.1.1. Hash-based vs Hashless
Hash-based algorithms exhibit space savings values close to each other across datasets. The best among the hashless algorithms (AE-Max, AE-Min, MAXP, and RAM) achieves slightly lower space savings than the best hash-based algorithm on some datasets, such as DEB and NEWS (Figures 8(a) and 8(f)). On the other hand, the best hashless algorithm outperforms all hash-based algorithms on other datasets, such as LNX and RDS (Figures 8(e) and 8(g)). Overall, the best hashless algorithms achieve space savings values within 6% of the best hash-based ones across all datasets and chunk sizes.
The only exception to this is MAPS (Figure 8(j)), where hashless algorithms perform worse than hash-based ones. AE-Max, the best performing hashless algorithm, achieves 11% less space savings than CRC, its best hash-based counterpart. This suggests that there are some specific datasets in which hash-based algorithms are the optimal choice. However, in most cases, hashless algorithms remain competitive with their hash-based counterparts for data deduplication.
6.1.2. Hashless algorithm comparison
The best-performing hashless algorithm depends on the dataset’s characteristics and the average chunk size. For instance, RAM achieves the highest space savings on DEB (Figure 8(a)) while MAXP does so on TPCC (Figure 8(h)). This shows that accelerating all hashless algorithms is important, as the optimal algorithm depends on the dataset’s characteristics.
(a) AE-Max / VAE-Max
(b) AE-Min / VAE-Min
(c) MAXP / VMAXP
(d) RAM / VRAM
Figure 9. Chunk size distribution CDFs of hashless algorithms and their AVX-512 accelerated versions on TPCC with an 8KB average chunk size
AE-Max either performs equivalently to or outperforms AE-Min on all datasets, except DEB and TPCC (Figures 8(a) and 8(h)). In addition, note that AE-Min is adversely affected by the byte-shifting pattern in MAPS, causing it to achieve only 8.89% in space savings while other CDC algorithms achieve 58%-78%.
Finally, while MAXP achieves higher space savings than RAM and both AE variants on many datasets, the space savings difference between it and the next best hashless algorithm is small.
6.1.3. Vector-acceleration Impact
Note that SS-CRC, SS-Gear, VAE, VMAXP, and VRAM achieve the same space savings as their native counterparts, i.e., vector-acceleration does not impact the space savings achieved by CDC algorithms. This aligns with the results previously observed for SS-CRC and SS-GEAR (sscdc, ). We omit these results from Figure 8 for clarity.
6.1.4. Chunk Size Distribution
To ensure the correctness of our vector-accelerated implementations, we compared their chunk size distributions to those of their native counterparts and verified that they were equal. Figure 9 shows the chunk size distributions exhibited by AE-Max, AE-Min, MAXP, and RAM compared against their AVX-512 versions accelerated with VectorCDC. Note that each figure is a cumulative frequency (CDF) (cumulativefreq, ) plot. We use a target average chunk size of 8KB and the TPCC dataset for this experiment. The results for other datasets and chunk sizes were similar and have been omitted for clarity.
Figure 9 shows that the AVX-512 versions of each algorithm exhibit the same chunk size distribution as their unaccelerated counterparts. For instance, Figure 9(d) shows that RAM and VRAM have identical chunk size distributions. Thus, VectorCDC does not affect the chunks generated by the hashless algorithms it accelerates.
6.2. Chunking Throughput
(a) DEB
(b) DEV
(c) Hash-based algorithms with SS-CDC (sscdc, )
(d) Hashless algorithms with VectorCDC
Figure 10. Chunking Throughput with AVX-512 instructions and 8KB chunks. Note that the legend entries are in the same order as the plot bars from Figures 10(a) and 10(b).
Figures 10(a) and 10(b) show the throughput achieved by all algorithms on DEB and DEV with a chunk size of 8KB. Note that vector-accelerated algorithms are shown with patterned bars and that we have cropped the y-axis to 5 GB/s to avoid the figures being skewed by VRAM. The results on other datasets and chunk sizes had similar trends and have been omitted for clarity.
6.2.1. Throughput Comparison
Figures 10(a) and 10(b) show that hashless algorithms accelerated with VectorCDC achieve the highest throughputs among all CDC algorithms. VRAM, the fastest accelerated hashless algorithm, achieves $8.35\times$ and $15.3\times$ higher throughput than SS-GEAR and FastCDC, the fastest accelerated and unaccelerated hash-based algorithms, respectively.
Among unaccelerated hash-based algorithms, FastCDC (fastcdc, ), Gear (gear_hash, ), and CRC (sscdc, ) are the fastest, achieving 1.95 GB/s, 1.05 GB/s, and 0.48 GB/s, respectively. We accelerated each of these using SS-CDC (sscdc, ); SS-GEAR achieves 3.57 GB/s and SS-CRC achieves 1.17 GB/s. We discuss accelerating FastCDC in §6.2.2.
The unaccelerated hashless algorithms AE-Max (ae, ), AE-Min (ae, ), MAXP (maxp, ), and RAM (ram, ) come in at $1.5$ GB/s–$1.7$ GB/s, faster than most of their hash-based counterparts. Their VectorCDC-accelerated versions, VAE-Max, VAE-Min, VMAXP, and VRAM, achieve $6.5$ GB/s–$29.9$ GB/s. VRAM, the fastest accelerated hashless algorithm, is $8.35\times$ and $26.2\times$ faster than SS-GEAR and SS-CRC respectively. While the throughputs of all algorithms slightly vary across datasets, the trends remain the same.
6.2.2. Accelerating FastCDC.
We did not observe any speedup when accelerating FastCDC (fastcdc, ) with SS-CDC (sscdc, ). One of the main throughput optimizations used by FastCDC is sub-minimum skipping (§2.1). However, as noted in §3.2, decoupling the rolling-hash phase from the boundary identification phase eliminates the throughput benefits of minimum chunk size skipping, nullifying any speedup provided by vector-acceleration. VAE-Max, VAE-Min, VMAXP and VRAM achieve $6\times, 4.4$$46\times$, $4.76\times$, and $15.3\times$ higher throughput than FastCDC respectively.
6.2.3. Vector-acceleration benefits.
Figures 10(c) and 10(d) compare the throughput benefits of accelerating hash-based and hashless algorithms with AVX-512 accelerated algorithms on DEB and DEV.
Accelerating hash-based algorithms (Figure 10(c)) using SS-CDC achieves a speedup of $2.45-3.32\times$. For instance, on DEB, SS-CRC exhibits a throughput of $1.2$ GB/s over CRC at $0.46$ GB/s, i.e., a $2.45\times$ speedup. Similarly, SS-GEAR achieves $3.57$ GB/s over GEAR at $1.07$ GB/s, i.e., a $3.32\times$ speedup.
On the other hand, the hashless algorithms VAE-Max, VAE-Min, VMAXP, and VRAM achieve speedups of $5.1\times$, $4.43\times$, $5.36\times$, and $17.69\times$ over their respective native counterparts. Thus, vector instructions can be leveraged far more efficiently for hashless algorithms, proving that hashless algorithms are better candidates for vector-acceleration than their hash-based counterparts.
Figure 10(d) shows that VRAM achieves higher throughputs than VAE-Max, VAE-Min, and VMAXP. This is because VAE requires multiple iterations of Range Scan per chunk, each followed by an Extreme Byte Search, while VRAM only requires one iteration of each (§4). Similarly, VMAXP requires multiple Range Scans, each followed by two Extreme Byte Searches. For a given target average chunk size, the size of the Extreme Byte Search regions in MAXP is 708070-80% smaller than the search region in AE. This allows VMAXP to achieve speeds similar to VAE-Max and VAE-Min despite needing an extra Extreme Byte Search.
Thus, RAM is inherently more vector-friendly than AE and MAXP. However, note that VAE and VMAXP are still faster than every other CDC algorithm.
6.2.4. Deduplication performance bottlenecks.
Figure 11 shows the time taken by the chunking and hashing phases in the deduplication pipeline on DEB with an 8KB average chunk size. We omit the results for other datasets as they were similar. We used xxHash3, the fastest hashing algorithm (§3.1) for fingerprinting and ran this experiment on the Intel Emerald Rapids machine. We use AVX-512 versions of hashless CDC algorithms, accelerated with VectorCDC. Note that the times in Figure 3 and Figure 11 differ due to the differing sizes and dataset characteristics between the random dataset used in §3 and DEB.
Figure 11. Time taken for Data Chunking vs Fingerprinting on DEB with an 8 KB average chunk size, xxHash3 for fingerprinting, and AVX-512 instructions for acceleration.
For unaccelerated CDC algorithms, Figure 11 shows similar results to those obtained in Figure 3 with random data. Data chunking takes significantly longer than fingerprinting with these algorithms. On the other hand, VAE-Min, VAE-Max, VMAXP, and VRAM show data chunking times similar to or lower than fingerprinting. For instance, with VRAM, data chunking takes $1.29$ seconds while fingerprinting takes $2.27$ seconds.
This shows that VectorCDC effectively alleviates the data chunking bottleneck in the deduplication pipeline.
6.3. Throughput breakdown - Extreme Byte Search vs Range Scan
(a) VRAM
(b) VMAXP
Figure 12. Throughput Breakdown with AVX-512 instructions
(a) DEB
(b) LNX
Figure 13. Percentage share of bytes processed using Extreme Byte Search and Range Scan by hashless CDC algorithms on DEB and LNX
The throughput impact of each processing pattern depends on algorithmic and dataset characteristics. Figure 12 shows the individual impact of accelerating Extreme Byte Search and Range Scan in VRAM on DEB and LNX with an 8KB chunk size. VRAM-EBS represents RAM running with only Extreme Byte Search acceleration, while VRAM-512 uses both accelerated patterns. Similarly, VMAXP-EBS represents MAXP running with only Extreme Byte Search acceleration, while VMAXP-512 uses both accelerated patterns. We use the DEB and LNX datasets for this experiment.
Figure 12(a) shows that on DEB, VRAM-EBS achieves a throughput of $18.5$ GB/s compared to RAM at $1.7$ GB/s. Accelerating Range Scan provides an additional speedup of $11.4$ GB/s. On the other hand on LNX, VRAM-EBS only achieves $2.7$ GB/s compared to RAM at $2$ GB/s/. Accelerating Range Scan provides an additional speedup of $27.6$ GB/s. Thus, each pattern has a balanced impact on VRAM’s throughput on DEB while Range Scan majorly contributes to throughput on LNX, showing that dataset characteristics impact the throughput breakdown.
The throughput breakdown also varies across algorithms; for instance, accelerating Extreme Byte Searches has differing impacts on the throughputs of RAM and MAXP. While Figure 12(a) shows that VRAM-EBS achieves significantly higher throughput than RAM on DEB, Figure 12(b) shows that VMAXP-EBS only achieves small speedups over MAXP, i.e., Range Scan acceleration contributes more to throughput on VMAXP than it does on VRAM.
These results are directly tied to the number of bytes processed by the algorithms on both datasets. Figure 13 shows the percentage shares of bytes processed by Extreme Byte Searches and Range Scans, for all hashless algorithms on DEB and LNX. As seen in Figure 13(a), the percentage shares differ across algorithms. For instance, RAM processes $96.70\%$ and $3.30\%$ of bytes on DEB with Extreme Byte Search and Range Scan, respectively. On the other hand, MAXP processes $10.26\%$ and $89.74\%$ of bytes with Extreme Byte Search and Range Scan, respectively. Additionally, this percentage varies across datasets, as seen by the differences between Figures 13(a) and 13(b).
Thus, accelerating both phases using vector instructions is crucial to performance, as the impact of each phase depends on dataset and algorithmic characteristics.
6.4. VectorCDC across different vector instruction sets
SS-CDC (sscdc, ) requires CPUs with vector instruction sets supporting scatter/gather instructions (§3.2). Such CPUs are present only in a small percentage of datacenter nodes today. However, almost all datacenter machines that are currently deployed support other vector instruction sets without this capability, such as SSE-128, AVX-256, NEON-128, and VSX-128 (§2.3).
(a) AMD EPYC Rome
(b) Intel Emerald Rapids
(c) Intel Skylake
(d) ARM v8 Atlas
(e) IBM Power 8
Figure 14. Accelerating hashless algorithms with VectorCDC across processor architectures, on DEB at an 8KB average chunk size. Note the different y-axis scale on Figures 14(d) and 14(e).
While §4 discusses VectorCDC’s design using AVX-512 instructions, the same methods can be applied to any vector instruction set that supports VCMP, VMAX, and VMASK operations. In this section, we evaluate VectorCDC’s performance with other such vector instruction sets. Figure 14 shows the throughput achieved by AE-Max, AE-Min, MAXP, and RAM accelerated using VectorCDC with SSE-128, AVX-256, NEON-128, and VSX-128 instruction sets. We ran this experiment using the DEB dataset and an average chunk size of 8 KB.
In summary, VectorCDC is compatible with a large range of platforms that support vector instructions while retaining its benefits over hash-based algorithms, unlike SS-CDC (sscdc, ) which requires CPUs with scatter/gather instruction support. While VectorCDC’s speedups depend on the underlying CPU’s vector instruction support, it remains effective on platforms alternative to x86, such as ARM and IBM Power.
6.4.1. AMD EPYC Rome
Figure 14(a) shows the throughputs achieved by hashless algorithms accelerated with VectorCDC on an AMD EPYC Rome machine. As shown in Table 1, the AMD machine only supports SSE-128 and AVX-256 instructions. All four hashless algorithms in Figure 14(a) show speedups over their native versions with both instruction sets. For instance, AE-Max achieves $2.12\times$ and $3.43\times$ speedups with SSE-128 and AVX-256 instructions, respectively. Similar to the results in §6.2 with AVX-512 instructions, RAM achieves the highest throughput of all algorithms with both SSE-128 and AVX-256 instructions.
6.4.2. Intel Emerald Rapids and Skylake
Figures 14(b) and 14(c) show the throughputs achieved by hashless algorithms accelerated with VectorCDC on Intel Emerald Rapids and Skylake machines. As shown in Table 1, these machines support SSE-128, AVX-256, and AVX-512 instructions. All four hashless algorithms in Figures 14(b) and 14(c) achieve speedups over their unaccelerated versions with all instruction sets. For instance, in Figure 14(c), AE-Max achieves $2.29\times$, $4.91\times$, and $6.71\times$ speedups with SSE-128, AVX-256, and AVX-512 instructions, respectively. Similar to the results in §6.2 with AVX-512 instructions, RAM achieves the highest throughput of all algorithms with both SSE-128 and AVX-256 instructions.
On both platforms, all algorithms also benefit from increasing vector widths; that is, higher vector widths lead to higher throughput. The only exception is MAXP, which does not gain as much as the other algorithms with AVX-512 instructions over AVX-256. This is related to the small window sizes used by MAXP for its Extreme Byte Search phases, which do not benefit from high vector widths. However, MAXP still achieves $4.7\times$ and $5.39\times$ speedups with AVX-512 instructions over its unaccelerated version, on the Skylake and Emerald Rapids machines, respectively.
6.4.3. ARM v8 Atlas
Figure 14(d) shows the throughputs achieved by hashless algorithms accelerated with VectorCDC on an ARM v8 Atlas machine. As shown in Table 1, the machine only supports NEON-128 instructions, an ARM equivalent to SSE-128. While the instruction set supports VMAX and VCMP operations, it lacks native support for VMASK operations (§5). RAM achieves the highest throughput among all accelerated hashless algorithms at $2.91$ GB/s.
All hashless algorithms achieve lower speedups on ARM with NEON-128 instructions, when compared to SSE-128 instructions on Intel and AMD machines. AE-Max and AE-Min are especially affected, achieving only $1.08\times$ and $1.05\times$ speedups, i.e. 8% and 5% gains with NEON-128 over their unaccelerated versions. This is largely due to the lack of native VMASK support, which affects Range Scans. While our implementation uses an alternative method to achieve the same functionality, it uses four NEON-128 instructions instead of a single SSE-128 VMASK instruction.
However, MAXP and RAM still achieve $1.93\times$ and $5.32\times$ speedups, respectively, showing that VectorCDC remains beneficial on ARM platforms with NEON-128 support. Note that these numbers are expected to improve in ARM platforms supporting SVE/SVE2 instructions (arm_sve, ), as they offer native VMASK support. However, as we could not obtain such a platform for our evaluation, we leave a detailed SVE/SVE2 performance review as future work.
6.4.4. IBM Power 8
Figure 14(e) shows the throughputs achieved by hashless algorithms accelerated with VectorCDC on an IBM Power 8 machine. As shown in Table 1, this machine only supports VSX-128 instructions, an IBM equivalent to SSE-128. This instruction set lacks support for native VMASK operations as well (§5). RAM achieves the highest throughput among all accelerated hashless algorithms, at $8.54$ GB/s.
Unlike ARM, all hashless algorithms exhibit considerable speedups on IBM Power 8 with VectorCDC. AE-Max and AE-Min achieve speedups of $2.92\times$ and $2.85\times$ respectively. MAXP and RAM achieve speedups of $7.93\times$ and $20.35\times$ respectively. Furthermore, all hashless algorithms accelerated with VSX-128 instructions achieve speedups equivalent to or greater than their counterparts accelerated with SSE-128 on Intel and AMD machines. For instance, RAM achieves a speedup of $20.35\times$ with VSX-128 on IBM Power 8 while it achieves a speedup of $7.49\times$ and $9.94\times$ with SSE-128 on Intel Emerald Rapids and AMD EPYC Rome, respectively.
This is because, despite the lack of native VMASK instruction support, the alternative implementation using vec_bpermq is efficient and uses just two fast VSX-128 instructions.
7. Related Work
7.0.1. Chunking optimizations.
Many efforts have been made to optimize data chunking. MUCH (much, ) and P-Dedupe (p_dedupe, ) use multiple threads to accelerate chunking. RapidCDC (rapidcdc, ) sometimes skips data chunking by predicting the next chunk boundary based on historical data, but requires maintaining additional metadata. Bimodal Chunking (kruus2010bimodal, ) initially splits the data into large chunks, and then divides duplicate adjacent chunks into smaller ones, to enhance space savings. VectorCDC is compatible with all of these approaches, as they build on top of existing CDC algorithms.
Previous work (lowentropy_cloud, ) which analyzes the characteristics of chunks generated by CDC algorithms, is orthogonal to VectorCDC, as vector acceleration does not affect generated chunks.
7.0.2. Deduplication optimizations.
Several other efforts exist to optimize the other phases of the deduplication pipeline. StoreGPU (storegpu, ) and GPU-Dedup (gpu_dedup, ) accelerate chunk hash computation using GPUs. SiLo (Silo, ), Sparse Indexing (sparse_indexing, ) and Extreme Binning (extremeBinning, ) optimize hash indexing. HYDRAStor (hydrastor, ) is a distributed deduplication system that focuses on data placement. Several studies incorporate delta compression after deduplication to further compress similar but non-duplicate chunks (shilane2012wan, ; zou2022building, ; zhang2023loopdelta, ). These efforts are orthogonal to ours as we accelerate the data chunking phase.
7.0.3. Accelerating other storage systems.
Vector instructions have been widely used to accelerate other storage systems. MinervaFS (minervafs, ) accelerates the computation of transform and basis functions in generalized deduplication with AVX instructions. ICID (inlinededup_nvme, ) records memory-copy operations in a B-Tree for fine-grained deduplication, accelerating tree searches with AVX instructions. AVX-512 conflict detection instructions have been used to accelerate lightweight data compression algorithms (avx_rle, ). Numerous works attempt to accelerate collision-resistant hashing algorithms used across storage systems with vector instructions (clhash_vectorized, ; murmur_vectorized, ). These efforts are orthogonal to ours as we focus on using vector instructions to accelerate CDC algorithms for block-level deduplication.
7.0.4. Secure deduplication systems.
Several efforts build end-to-end deduplication systems for encrypted data (secure_dedup_survey, ). They mainly target encryption schemes (mle, ; secure_asokan, ) for the underlying data or focus on reducing attacks on the system (side_channel_dedup, ; sgxdedup, ). Some target specific applications, such as distributing encrypted docker images  (sun2024simenc, ) and encrypted videos (securededup_video, ). As all of these efforts layer encryption atop existing data chunking algorithms, VectorCDC is compatible with all these approaches.
8. Conclusion
We present VectorCDC, a methodology for accelerating content-defined chunking using vector instructions. VectorCDC avoids the pitfalls of previous work that accelerates CDC algorithms by choosing hashless CDC instead. VectorCDC accelerates these algorithms using novel tree-based search and packed scanning methods. Our evaluation shows that VectorCDC achieves $8.35\times$-$26.2\times$ higher throughput than existing AVX-based CDC techniques. We have made our code publicly available by integrating it with DedupBench (dedupbench, ), and hosted one of our datasets on Kaggle (deb_dataset, ).
Acknowledgements.
We thank the anonymous USENIX FAST 2025 and ACM Transactions on Storage reviewers for their feedback. We thank Lori Paniak for his technical assistance throughout the project, and Mu’men Al-Jarah for his feedback on an earlier version of this work. The research team was supported by grants from the National Cybersecurity Consortium (NCC), Natural Sciences and Engineering Research Council of Canada (NSERC), and the Ontario Research Fund’s Research Excellence Program (ALLRP-561423-20, RGPIN-2025-03332, and ORF-RE012-051). The team was also supported by research grants from Acronis, Oracle Research Labs, and Rogers Communications. Sreeharsha is supported by the Cheriton Graduate Scholarship and the Ontario Graduate Scholarship.